A tea mixture was made from 30lb of tea that cost $6.00 per pound and 70lb of tea that cost $3.20 per poumd.Find the cost per pound of the tea mixture?

Answer by Maths68(1474)

Tea Type A
Price per lb = $6.00
Cost of 30 lb = 30*6=$180
=================================
Tea Type B
Price per lb = $3.20
Cost of 70 lb = 70*3.20=$224
=================================
Mixture
Type A + Type B = 30+70 = 100 pound
Price of mixture = Price of Type A + Price of Type B
Price of mixture = 180+224 = $404
=============================================
Price per pound of mixture = Total price of mixture/Total amount of mixture=404/100=$4.04/pound